OpenStudy (anonymous):

Mike has $1.55 in dimes and nickels. If he has 7 more nickels than dimes, how many of each coin does he have?

OpenStudy (anonymous):

@Luigi0210 close this

OpenStudy (anonymous):

i thought i knew the answer to this quest but i am screwed this morning...sorry

OpenStudy (anonymous):

That's okay. Thank you for trying. <3

OpenStudy (anonymous):

wats 35/64 i need help

OpenStudy (anonymous):

0.546875

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Take no. of nickels as x and no. of dimes as y dime=10 cent nickel = 5 cent So 5x + 10y =155 cents x = y + 7 Solving 5( y + 7) + 10y = 155 y = 8 x = 15

OpenStudy (anonymous):

Thank you very much. ^w^

OpenStudy (anonymous):

No problem